OpenStudy (anonymous):

evaluate integral from o to pi 36+4cos^2(theta) d(theta)

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi}1/2(36+4\cos^2)\theta\]

OpenStudy (anonymous):

that is the one